OpenStudy (anonymous):

Calculus: The Comparison Test I'm supposed to show if convergent or divergent using the comparison test or limit comparison test: (n-1)!/(n+2)!

OpenStudy (anonymous):

use limit comparison test and then u can show that this series is convergent.

OpenStudy (anonymous):

u know limit comparison test?

OpenStudy (anonymous):

or u can show using comparison test also.

OpenStudy (anonymous):

(n-1)!/(n+2)! = 1/n(n+1)(n+2) ---a(n) look for some other series---b(n) about which u know that either it's convergent or divergent suppose b(n)= 1/n^3 and this is convergent . a(n) and b(n) are +ve for all n. a(n)< b(n) since b(n) is convergent so a(n) will also converge.

OpenStudy (anonymous):

does it help u?

OpenStudy (anonymous):

Yes! Thanks!

OpenStudy (anonymous):

:)